assume temp. remains const. in uppe part of atmosphere.obtain expression for variation in pressure in upper atmosphere with height.the mean molecular weight of air is M. acceleration due to gravity is g.

Grade:12

2 Answers

shruti gupta
14 years ago
Hi Shefali, if for a small height diff. of (dh) the change in pressure is (dp) then dp = Dgdh (d is the density of the gas at height h) D=PM/RT (by gas laws) dp=-PMgdh/RT (-ve sign indicates decrease in pressure with height) dp/dh= -PMgdh/RT I hope it will clarify your doubt
shruti gupta
14 years ago
hi shefali if you consider the change in height dh for which the pressure change is dp then dp=-Dgdh Disthe density of air at height h ,-ve sign indicates decrease in pressure with increase in height D=PM/RT dp=-PMgdh/RT dp/dh=-PMg/RT i hope it will clarify your doubt.